VEGAS Team Meeting

February 11, 2013

Attendees

CALTECH, UC-B, & Columbia

Dan, Jayanth, Mark, Andrew, Glenn

NRAO:

Jason, Ray, Marty, Anish, Randy, Paul, John

Meeting Notes:

Next Teleconference: February 18, 2012 .

Postpone observations until the receiver is replaced.

Latest test results

Action Items from Last meeting
  1. Ray to make VEGAS simulator.
    • (11-Feb)
  2. A BOF for mode 13, or some other mode that uses GPUs.
    • (04-Feb) Not available yet
    • (11-Feb) 16 input version compiling now for subsequent test
  3. A BOF with working switching signal logic.
    • (04-Feb) designed but still troubleshooting switching table loading issues
    • (11-Feb) Added to design for Mode #1
    • (11-Feb) Add sfp shift into manager
  4. A test and decision on what computers to buy.
    • (04-Feb) Not yet determined - Jayanth should go ahead with his tests to complete selection
    • (04-Feb) Needs HPC11.
    • (11-Feb)
  5. All Banks not starting at the same time.
    • (04-Feb) (above)
    • (11-Feb)
  6. It will be good if we get the full spectral dynamic range corresponding to the 8 bit ADC in the next version of mode 1.
    • (04-Feb) Conduct initial tests at Berkeley, to see if there are any limiting issues for dynamic range and optimize
    • (04-Feb) what is/are the requirement(s) w.r.t full scale sin wave?
    • (04-Feb) is this a new requirement for the manager (yes)? is a parameter required to select between options? (parametrized shifting)
    • (04-Feb) what are requirements for shift parameter? depends on result of studying spectra...
    • (11-Feb)
  7. KFPA observation with mode 13.
    • (04-Feb) awaits mode 13
    • (11-Feb)
  8. Manager/VEGAS HPC code that will let two instances of the program be run simultaneously.
    • (04-Feb) not available, but not required for GC obs.
    • (11-Feb)
  9. Remember to do the RFI test of the rack.
    • (04-Feb) Not yet scheduled
    • (11-Feb)
  10. Roach to Valon serial port connection.
    • (04-Feb) Not yet implemented; not req'd for GC obs
    • (11-Feb)
  11. The I2C bus software tested with the manager and Galen's modules.
    • (04-Feb) Post GC Obs
    • (11-Feb)
  12. The ROACH uboots reflased (1-6).
    • (04-Feb) May be accomplished any time with access to system (Feb 5)
    • (11-Feb) Complete
  13. The tcpborphserver init script needs configured.
    • (04-Feb) not done, not req'd for GC obs
    • (11-Feb) Complete
  14. Add sftware version onformation into FITS file
    • (11-Feb) After GC obs

VEGAS observation preparation Status:

  1. Long Artificial Pulsar Test (Ray + Paul starting 11-Feb)
    • Longer (3-hrs) x2
    • Ray to look at start up issues and then many ~5 min scans
  2. Request Ku-Band receiver installation
    • If long delay for installation, K-Band Pulsar test (reduced weather requirements)

Action Items:

Other items:

-- MartyBloss - 2013-02-11
Topic revision: r2 - 2013-02-11, MartyBloss
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback